OpenStudy (anonymous):

@mathman806 The dimensions of sphere B are 3 times the dimensions of sphere A. If the volume of sphere A is 26 in3, what is the volume of sphere B?

OpenStudy (anonymous):

well sphere b = sphere a so b=3a substitute in 26 for a b=26*3 what do you get?

OpenStudy (anonymous):

78??

OpenStudy (anonymous):

i was 702

OpenStudy (anonymous):

78in^3

OpenStudy (anonymous):

what about this one?? The dimensions of cube B are 3 times the dimensions of cube A. If the surface area of cube A is 6 cm2, what is the surface area of cube B?

OpenStudy (anonymous):

do the same thing as the last one b=3a a=6 b=3*6 b=18

OpenStudy (anonymous):

thats not an answer

OpenStudy (anonymous):

um what are the choices?

OpenStudy (anonymous):

54 cm2 24 cm2 48 cm2 26 cm2

OpenStudy (anonymous):

then idk

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Rectangles ABCD and EFGH are similar figures because angles A, B, C, and D are congruent to angles E, F, G, and H, respectively. If DA equals 15 units, AB equals 10 units, and HE equals 75 units, what does EF equal?

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

first off make a ratio 75:15 reduce 5:1 now plug in the ratio into an equation 5x=y if ab is similar ef then 5*10=y multiply 50=y EF=50

OpenStudy (anonymous):

Right triangle ABC is similar to right triangle DEF. If the measure of angle A is 36°, what is the measure of angle F?

OpenStudy (anonymous):

angle a = angle d angle b = angle e angle c = angle f angles b & e are 90 degrees b-a=c e-d=f 90-36=c 90-36=f solve

OpenStudy (anonymous):

54...???

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

The shadow of a 4-foot tall post is 16 feet long. At the same time a tree casts a shadow that is 44 feet long. What is the height of the tree?

OpenStudy (anonymous):

@mathman806 please i really need to pass this

OpenStudy (anonymous):

sorry been off for a long time yes 54 is right for the second one 4/16=x/44 cross multiply 4*44=16x multiply 16x=176 divide by 16 x=11